If the son is aged 5 and his father is 30 when will his father be voluble the sons age ?
Answer by ewatrrr(24785)



Hi
30+x = 2(5 + x)
30 + x = 10 + 2x
20 = x
20 years from now.
Checking...
50 = 2*25